OpenStudy (anonymous):

If final velocity is 0 , initial velocity is positive and acceleration is positive , wouldn't time be negative @mathslover @Miracrown

mathslover (mathslover):

Acceleration can't be positive in that case.

mathslover (mathslover):

As the velocity is \(\bf{decreasing}\) so, there will be deceleration .

OpenStudy (anonymous):

Kinematic equations do not work when there are other external factors

OpenStudy (anonymous):

?

OpenStudy (anonymous):

r u there @mathslover

mathslover (mathslover):

give me some time, working on a question

OpenStudy (anonymous):

sure

mathslover (mathslover):

Yeah, Mira is here. She will help you :-)

Miracrown (miracrown):

Sure. Could you be more explicit? @No.name ?

Miracrown (miracrown):

is it about a UNIFORMLY ACCELERATED motion in one dimension?

mathslover (mathslover):

v - u is negative here a = v - u / t t = v - u /a t is always positive... t = v - u / a = - ve /a = +ve a = -ve ( the direction of a will be opposite to the direction of motion of the object) this is what I think.

mathslover (mathslover):

Conditions applied ^

OpenStudy (anonymous):

uhh , yeah uniform acceleration motion in 1d

Miracrown (miracrown):

could you share the original text with me?

mathslover (mathslover):

I'm getting used to how @Miracrown doesn't use LaTeX and answers the questions very well... probably, I will start writing like her soon! =P I forgot LaTeX !!!!!!!!!!

OpenStudy (anonymous):

This is my own query

Miracrown (miracrown):

Yeah me no need of latex, :P

OpenStudy (anonymous):